OpenStudy (anonymous):

Find the value of the expression below. 2 6 8

OpenStudy (mrnood):

are you going to give us the expression?

OpenStudy (anonymous):

it doesnt have one

OpenStudy (anonymous):

yeeeeeah jjmack w cant solve the question when there is none

OpenStudy (mrnood):

is this meant to be \[2\div6\div8\]

OpenStudy (anonymous):

ig 2 6 and 8 are one of the expressions

OpenStudy (anonymous):

or 2*6*8

OpenStudy (anonymous):

2*6*8=96

OpenStudy (anonymous):

2/6/8=0.0416666667

OpenStudy (mrnood):

you need to be more clear about the question what you have written is not an expression - so we cannot help you to evaluate it

OpenStudy (anonymous):

or if it is 2+6+8=16

OpenStudy (mrnood):

don't confuse her further

OpenStudy (anonymous):

can i get a medal?

OpenStudy (anonymous):

i dont get it...........